OpenStudy (anonymous):

A stage is 30 feet long and has a perimeter of 110 feet. How wide is the stage?

OpenStudy (anonymous):

30+30=60 110-60=50 50/2=25 That means my answer is 25 ft.

OpenStudy (anonymous):

CLOSE QUESTION

OpenStudy (anonymous):

Yes, the answer is 25

OpenStudy (anonymous):

CLOSE QUESTION

OpenStudy (anonymous):